OpenStudy (princeevee):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? A. 9,40,42 B. 5,10,15 C. 6,8,10 D. 13,35,37 (((will medal)))

OpenStudy (mrnood):

out of each 3 numbers take the largest and square it then take the Other and squre them - add the 2 squared numbers together IF the sum of the 2 samller squares = the larger sqare then it is a right angle. FOR INSTANCE 3,5, 4 the largest is 5 and 5^2 = 25 the other numbers are 3, and 4 3^2 + 4^2 = 9+16 = 25 so sides of 3,4 & 5 make a right angle triangle do that with your sets of numbers to find the one where this is true...

OpenStudy (princeevee):

so c?

OpenStudy (mrnood):

it is best if YOU do the calculation. If someone gives the answer you will not have any better understanding. Pythagoras theorem says in a right angle triangle with sides a, b , c  if c is the longest side then a^2 + b^2 = c^2